# Parameter estimation of range-migrating targets using OTFS signals from LEO satellites

### v2

@@ V-B Extension to multiple targets
$\bar{r}_{\rm min},\,\bar{r}_{\rm max}]$ Inspected range interval [20, 39] km
@@
VI Numerical analysis
To assess the system performance, we consider the parameters in Table I. The maximum variation of the target range during the OTFS frame is $\bar{v}_{\rm max}NT\approx 136$ m, corresponding to about $3.5$ times the range resolution $R_{\rm r}$; the initial range $d$ and the range-rate $v$ are randomly generated, while a Swerling I fluctuation model is considered for the amplitude $\alpha$. The information symbols $\{X_{\rm DD}[k,l]\}$ are independently drawn from a unit-energy $4$-QAM constellation. The entries of the observed vector $\bm{z}$ are computed according to Eqs. (11), (14), and (18). Instead, the radar receiver is designed according to the echo model proposed in Secs. III and IV for ideal and rectangular shaping filters, respectively, that relies upon the $B$-fold stop-and-go approximation. In the implementation of the proposed two-step estimator, we set $\Sigma_{\rm r}=R_{\rm r}/100$, $\Sigma_{\rm rr}=R_{\rm rr}/100$, $\tilde{G}_{\rm r}=401$, $\tilde{G}_{\rm rr}=401$, $K=4P$, and $\epsilon=1$.
For future reference, we define the signal-to-noise ratio (SNR) of the target as $\mathrm{SNR}=\mathrm{E}[|\alpha|^{2}]NM/\sigma_{\omega}^{2}$. Also, let $d$, $v$, and $\alpha$ be the true values of the initial range, range-rate, and amplitude of the target, and let $\hat{d}$, $\hat{v}$, and $\hat{\alpha}$ be the corresponding estimates; then, we define the root mean square error (RMSE) in the estimation of the initial range, the RMSE in the estimation of the range-rate, and the NRMSE in the estimation of the amplitude as $\mathrm{RMSE}_{\rm r}=(\mathrm{E}[|d-\hat{d}|^{2}])^{1/2}$, $\mathrm{RMSE}_{\rm rr}=(\mathrm{E}[|v-\hat{v}|^{2}])^{1/2}$, and $\mathrm{NRMSE}_{\rm\alpha}=(\mathrm{E}\left[|\alpha-\hat{\alpha}|^{2}\right]/\mathrm{E}\left[|\alpha|^{2}\right])^{1/2}$, respectively. Next, we discuss two examples with one and multiple targets, respectively.
VI-A Single target
@@
Figure 3: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $B=1,2,4,8,16,32,64,128$ for $\mathrm{SNR}=10,20,30$ dB, when a high- or low-speed target is present and ideal shaping filters are employed. Both the ML and the proposed estimator are considered for comparison.
@@
Figure 4: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $\mathrm{SNR}$ for $B=4,8,16$, when a high-speed target is present and the proposed estimator is employed. Both ideal and rectangular shaping filters are considered.
@@
Figure 5: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $\bar{v}_{\rm min}$, when rectangular filters are used, $\mathrm{SNR}=15$ dB, and $\bar{v}_{\rm max}=\bar{v}_{\rm min}+1$ km/s. The proposed estimator is implemented for $B=1,2,4,8,16$. For comparison, the performance of the estimators presented in [21] and [29] is also evaluated.
@@
Next, we consider the use of rectangular shaping filters. Fig. 4 shows the behavior of $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\rm\alpha}$ as functions of $\mathrm{SNR}$ when the proposed estimator is employed and $B=4,8,16$. For comparison, results obtained with ideal shaping filters are also shown. Notably, the more practical rectangular filters achieve estimation performance that closely approximates that of the ideal filters.
Finally, we compare the performance of the proposed estimator with the methods in [21] and [29], which do not explicitly account for range migration within the OTFS frame. To evaluate performance across varying mobility scenarios, we set $\bar{v}_{\rm max}=\bar{v}_{\rm min}+1$ km/s and vary $\bar{v}_{\rm min}$ from 0 to 15 km/s. Fig. 5 plots the metrics $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\alpha}$ versus $\bar{v}_{\rm min}$ for rectangular shaping filters and $\mathrm{SNR}=15$ dB. The performance of the estimators in [21] and [29] deteriorates rapidly when $\bar{v}_{\rm min}$ exceeds 1 km/s. In this regime, the maximum range variation during the OTFS frame, given by $\bar{v}_{\rm max}NT$, becomes comparable to or exceeds the range resolution $R_{\rm r}$, resulting in a mismatch between the assumed and actual echo models. In contrast, the proposed estimator maintains robust performance across a broader range of $\bar{v}_{\rm min}$ by increasing $B$, thereby enabling a finer stop-and-go approximation in (7). Performance degradation only occurs when the range variation over $N/B$ symbol intervals, i.e., $\bar{v}_{\rm max}NT/B$, becomes comparable to or exceeds $R_{\rm r}$. Notably, with $B=16$, the proposed estimator delivers consistent performance across all tested values of $\bar{v}_{\rm min}$, in keeping with the findings in Fig. 3. Furthermore, as $\bar{v}_{\rm min}\rightarrow 0$, its performance aligns with that of the existing methods, confirming its effectiveness and practical relevance even for low-speed targets.
VI-B Multiple targets
@@
Figure 7: $|\bm{F}[k,l]|/\sigma_{\omega}$ vs $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ at the first (top), second (middle), and third (bottom) iteration of the CLEAN procedure, when rectangular filters are employed, $P=3$ high-speed targets are present, $B=16$, $\mathrm{SNR}_{1}=30$ dB, $\mathrm{SNR}_{2}=20$ dB, and $\mathrm{SNR}_{3}=10$ dB.
@@
We use here rectangular shaping filters and consider $P=3$ targets. Let $\mathrm{SNR}_{p}$ be the SNR of Target $p$ for $p=1,\ldots,P$. For each target, Fig. 6 reports the metrics $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\rm\alpha}$ versus $\mathrm{SNR}_{2}$, when $B=16$, $\mathrm{SNR}_{1}/\mathrm{SNR}_{2}=\mathrm{SNR}_{2}/\mathrm{SNR}_{3}=10$, and the proposed CLEAN-based estimator is employed. For comparison, we also plot the achievable performance when only a single target is present and the ML estimator in (55) is employed: we refer to it as the single-target bound. The proposed procedure is able to effectively estimate the parameters of each target, providing a performance close to the single-target bound. For a single instance and $\mathrm{SNR}_{2}=20$ dB, Fig. 7 shows $|\tilde{F}[k,l]|/\sigma_{\omega}$ versus $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ in each iteration of the CLEAN procedure, where $\tilde{F}[k,l]$ is the target response sequence recovered by the BOMP algorithm. It is seen that the echo of Target 1 (the strongest one) is recovered in Iteration 1 and subtracted out from the observed data; then, the echoes of Targets 2 and 3 are recovered in Iterations 2 and 3, respectively.
VII Conclusions
@@
In this work, we have investigated the use of OTFS communication signals for estimating the parameters of space targets experiencing range migration, which may occur in LEO satellite applications. Our major technical contribution is the derivation of a novel model for the target echo, assuming that the target range changes $B$ times within an OTFS frame; in particular, we have shown that range migration causes a predictable spread in the sparse target response in the delay-Doppler domain. Additionally, we have developed an approximate implementation of the maximum likelihood estimator for the target’s initial range, range-rate, and amplitude, which avoids exhaustive search in the delay-Doppler domain. If combined with the CLEAN algorithm, the proposed procedure can also handle multiple targets. A significant finding is that the estimation accuracy for high-speed targets is similar to that for low-speed targets when range migration is properly accounted for at the design stage.
@@
Future studies should consider that the radar receiver may not have perfect knowledge of the emitted symbols. In such cases, the use of dedicated radar symbols may be considered, and the optimal resource allocation between sensing and communication tasks should be studied. Furthermore, using multiple coordinated LEO satellites for signal transmission and/or reception may allow for estimating the initial position and velocity vectors of the target rather than only its initial range and range-rate. Finally, the use of multiple OTFS frames of different sizes may provide additional degrees of freedom and help resolve possible range and range-rate ambiguities.
We provide here the proof to Proposition 1. Upon plugging (35) and (36) in (18), we have
@@
Upon expanding $\Pi_{k,l}^{\rm ici}[k^{\prime},l^{\prime}]$ as shown in (62) at the top of this page, we obtain
@@
$Y_{{\rm DD}}^{\rm ici}[k,l]$ $=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=\big\lceil M-\frac{r\left(bT_{B}\right)M}{cT}\big\rceil}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\mathrm{e}^{-\mathrm{i}2\pi\frac{k^{\prime}}{N}}$

### v3

@@ V Numerical analysis
$\bar{r}_{\rm min},\,\bar{r}_{\rm max}]$ Inspected initial-range interval [20, 39] km
@@
V-A Single target
@@
Figure 4: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $B=1,2,4,8,16,32,64,128$ for $\mathrm{SNR}=10,20,30$ dB, when a high- or low-speed target is present and ideal shaping filters are employed. Both the ML and the proposed two-step estimator are considered for comparison.
@@
Figure 5: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $\mathrm{SNR}$ for $B=4,8,16$, when a high-speed target is present, and the proposed two-step estimator is employed. Both ideal and rectangular shaping filters are considered.
@@
Figure 6: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $B=1,2,4,8,16,32,64,128$ for different values NN and $\Delta$, when a high-speed target is present, rectangular shaping filters are employed, $\mathrm{SNR}=30$ dB, and the proposed two-step estimator is employed.
Refer to caption Refer to caption Refer to caption
Figure 7: $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus $\bar{v}_{\rm min}$, when rectangular filters are used, $\mathrm{SNR}=15$ dB, and $\bar{v}_{\rm max}=\bar{v}_{\rm min}+1$ km/s. The proposed two-step estimator is implemented for $B=1,2,4,8,16$. For comparison, the threshold-based estimator in [17] and the SOMP-based estimator in [24] are also included.
We first examine the system performance with ideal shaping filters. We then consider rectangular shaping filters.
V-A1 Ideal shaping filters
Fig. 4 presents the metrics $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\rm\alpha}$ as functions of $B$ for $\mathrm{SNR}=10,20,30$ dB. Observe that the parameters of a high-speed target cannot be accurately recovered when range variation during the OTFS frame is neglected at the system design (i.e., $B=1$). As $B$ increases from 1 to 16, estimation accuracy improves significantly, while further increases in $B$ yield diminishing returns. This behavior is governed by the ratio between the maximum range variation during one block of $N/B$ symbols and the range resolution, namely
$\frac{\bar{v}_{\max}NT/B}{R_{\rm r}}=\frac{\bar{v}_{\max}MN}{cB},$ (61)
which should be sufficiently smaller than one for accurate modeling. For the considered parameters, this ratio equals approximately $3.5/B$. Hence, for $B=1$, the blockwise constant-range approximation is inaccurate, whereas for $B=16$ the ratio reduces to about $0.22$, corresponding to a maximum range variation over $N/B$ symbols of approximately $8.5$ m. Such variation can be neglected with little performance loss.
@@ V-A1 Ideal shaping filters
Finally, note that the proposed two-step estimator achieves performance close to that of the ML estimator across all scenarios considered in Fig. 4, at a much lower implementation complexity, underscoring its effectiveness and practical utility.
V-A2 Rectangular shaping filters
@@
Fig. 5 shows the behavior of $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\rm\alpha}$ as functions of $\mathrm{SNR}$ when the proposed two-step estimator is employed and $B=4,8,16$. For comparison, results obtained with ideal shaping filters are also shown. Notably, the more practical rectangular filters achieve performance close to that of the ideal filters.
Fig. 6 investigates the influence of the OTFS parameters $\Delta$ and $N$ on the estimation performance when the proposed two-step estimator is employed and $\mathrm{SNR}=30$ dB. The results show that the initial-range estimation accuracy improves as the subcarrier spacing $\Delta$ increases, since the occupied bandwidth $M\Delta$ becomes larger, whereas it is only marginally affected by variations of $N$ when SNR is kept constant. In contrast, the range-rate estimation accuracy benefits from increasing $N$ and reducing $\Delta$, since both choices enlarge the coherent observation time $NT$ (recall that $T=1/\Delta$). As for the impact of the parameter $B$, the ratio introduced in (61) remains the key quantity. Note that this ratio is independent of $\Delta$ and scales linearly with $N$. Therefore, the pairs $(N,\Delta)=(128,15\penalty 10000\ \text{kHz})$ and $(128,30\penalty 10000\ \text{kHz})$, represented by solid lines, exhibit a similar validity region for the blockwise constant-range approximation, whereas the remaining pairs with $N=256$, represented by dashed lines, require approximately twice the value of $B$ to achieve a comparable level of approximation accuracy.
Finally, we evaluate the estimation performance of the proposed two-step estimator across varying mobility scenarios; to this end, we set $\bar{v}_{\rm max}=\bar{v}_{\rm min}+1$ km/s and vary $\bar{v}_{\rm min}$ from 0 to 15 km/s. For comparison, we also include the threshold-based estimator in [17] and the SOMP-based estimator in [24], both of which are designed under the assumption of a constant integer range within the OTFS frame. To ensure a fair comparison, the input delay-Doppler sequence $X_{{\rm DD}}[k,l]$ contains a single non-zero entry located at $(k,l)=(64,256)$, whereas all remaining entries are set to zero to create a guard region. [^8] In this setup, the complexity of the proposed two-step estimator simplifies to $\mathcal{O}\big(MK^{2}N^{2}+K^{4}N^{3}+\tilde{G}_{\text{r}}\tilde{G}_{\text{rr}}MN\big)$ because the vector $\bm{x}$ contains a single non-zero entry at a known location and the matrix-vector multiplication in (53) becomes trivial during the second step. In contrast, the benchmarks from [17] and [24] both have complexity $\mathcal{O}(MN)$. Fig. 7 plots the metrics $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\alpha}$ versus $\bar{v}_{\rm min}$ for $\mathrm{SNR}=15$ dB. The performance of the estimators in [17] and [24] deteriorates rapidly when $\bar{v}_{\rm min}$ exceeds 1 km/s. In this regime, the maximum range variation during the OTFS frame, given by $\bar{v}_{\rm max}NT$, becomes comparable to or exceeds the range resolution $R_{\rm r}$, resulting in a mismatch between the assumed and actual echo models. In contrast, the proposed two-step estimator remains robust over a significantly wider mobility range by using larger values of $B$, which enable a finer blockwise constant-range approximation in (8). In particular, $B=16$ provides stable performance over all considered values of $\bar{v}_{\rm min}$, consistently with the trends already observed in Figs. 4 and 6. Even when $\bar{v}_{\rm min}\rightarrow 0$, the proposed estimator still outperforms the existing methods due to the second refinement step, which improves the estimation accuracy at the cost of higher computational complexity.
V-B Multiple targets
@@
Figure 9: $|\tilde{F}[k,l]|/\sigma_{\omega}$ vs $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ at the first (left), second (center), and third (right) iteration of the CLEAN procedure, when rectangular filters are employed, $P=3$ high-speed targets are present, $B=16$, $\mathrm{SNR}_{1}=30$ dB, $\mathrm{SNR}_{2}=20$ dB, and $\mathrm{SNR}_{3}=10$ dB.
Refer to caption Refer to caption Refer to caption
Figure 10: Average $\mathrm{RMSE}_{\rm r}$ (left), $\mathrm{RMSE}_{\rm rr}$ (center), and $\mathrm{NRMSE}_{\rm\alpha}$ (right) versus range-rate separation, when rectangular filters are used, $P=2$ high-speed targets with the same initial-range are present, $B=16$, and $\mathrm{SNR}_{1}=\mathrm{SNR}_{2}=30$ dB. The performance of the proposed CLEAN-based estimator and the corresponding single-target performance bound are reported.
@@
We use rectangular-shaped filters here and consider $P=3$ targets. Let $\mathrm{SNR}_{p}$ be the SNR of Target $p$ for $p=1,\ldots,P$. For each target, Fig. 8 reports the metrics $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\rm\alpha}$ versus $\mathrm{SNR}_{2}$, when $B=16$, $\mathrm{SNR}_{1}/\mathrm{SNR}_{2}=\mathrm{SNR}_{2}/\mathrm{SNR}_{3}=10$, and the proposed CLEAN-based estimator is employed. For comparison, we also plot the achievable performance when only a single target is present and the ML estimator in (52) is employed: we refer to it as the single-target bound. The proposed procedure effectively estimates the parameters of each target, achieving performance close to the single-target bound. For a single instance and $\mathrm{SNR}_{2}=20$ dB, Fig. 9 shows $|\tilde{F}[k,l]|/\sigma_{\omega}$ versus $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ in each iteration of the CLEAN procedure, where $\tilde{F}[k,l]$ is the sequence recovered by the BOMP algorithm. It is seen that the target echoes are iteratively recovered and subtracted out from the observed data.
Finally, we assess the ability of the proposed CLEAN-based estimator to resolve two targets having the same initial-range and SNR but different range-rates. Fig. 10 reports $\mathrm{RMSE}_{\rm r}$, $\mathrm{RMSE}_{\rm rr}$, and $\mathrm{NRMSE}_{\alpha}$ (averaged over the two targets) versus the range-rate separation. When the range-rate separation is below the nominal resolution $R_{\rm rr}$, the two targets are difficult to distinguish, and the estimation errors increase significantly. As the separation increases beyond $R_{\rm rr}$, the two targets become reliably separable, and the performance approaches the single-target bound. This confirms that $R_{\rm rr}$ provides a meaningful reference for range-rate discrimination.
VI Conclusions
@@
In this work, we have investigated the use of OTFS communication signals for estimating the parameters of space targets experiencing range migration, as may arise in LEO satellite applications. Our main technical contribution is the derivation of a novel echo model that allows the target range to vary $B$ times within one OTFS frame. In particular, we have shown that range migration induces a spreading of the target response while preserving its sparse structure in the delay-Doppler domain. We have also developed a two-step estimator for the target initial-range, range-rate, and amplitude, achieving performance close to that of the maximum likelihood estimator at substantially lower complexity. When combined with the CLEAN algorithm, the proposed procedure can handle multiple targets. A significant finding is that the achievable estimation accuracy for high-speed targets is comparable to that of low-speed targets when range migration is properly accounted for in the receiver design.
@@
Future studies should consider that the radar receiver may not have perfect knowledge of the emitted communication symbols. In such cases, dedicated radar symbols may be considered, and the resource allocation between sensing and communication tasks should be investigated. Furthermore, using multiple coordinated LEO satellites for signal transmission and reception may allow for estimating the initial position and velocity vectors of the target rather than only its initial-range and range-rate. Finally, the use of multiple OTFS frames of different sizes may provide additional degrees of freedom and help resolve possible initial-range and range-rate ambiguities.
Appendix A Proof of Proposition 1
Upon plugging (12) into (15), we have
which completes the proof.
Appendix B Proof of Proposition 2
After exploiting (3) and (20), Eq. (19) becomes
Finally, notice that we have
$=NM\alpha\Phi[k-k^{\prime},l-l^{\prime}],$
where (63) follows from (8).
Appendix C Proof of Proposition 3
Upon plugging (33) and (34) in (19), we have
@@ Appendix C Proof of Proposition 3
Upon expanding $\Pi_{k,l}^{\rm ici}[k^{\prime},l^{\prime}]$ as shown in (67) at the top of the next page, we obtain
@@
$Y_{{\rm DD}}^{\rm isi}[k,l]$ $=\frac{\alpha}{B}\sum_{b=0}^{B-1}\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=\big\lceil M-\frac{r\left(bT_{B}\right)M}{cT}\big\rceil}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\mathrm{e}^{-\mathrm{i}2\pi\frac{k^{\prime}}{N}}$